Introduction

This is the manual of a macropackage I wrote for printing the covers for cassettes. It’s quite flexible and quite easy to learn. I tried to keep this manual short, so it won’t cost you that much to print it. – Stijn Raaijmakers.

General form of a tape-file

    \input ⟨size-file⟩.cas
    \input cascover

\language⟨language⟩*

    \begintape
      \begincontents

\beginside ⟨Side⟩: ⟨Title⟩ - ⟨Author⟩ (⟨Dolby⟩) \\⟨Title⟩ - ⟨Author⟩, ⟨Time⟩*

⋮ \title ⟨Title⟩ - ⟨Author⟩*

\\⟨Title⟩ - ⟨Author⟩, ⟨Time⟩*

⋮

        \endside

\beginside ⟨Side⟩: ⟨Title⟩ - ⟨Author⟩ (⟨Dolby⟩)

⋮

        \endside
      \endcontents
      \beginfront

\fronttitle ⟨Title⟩ - ⟨Author⟩*

⋮

      \endfront
      \beginflap

⋮

      \endflap
    \endtape
    \begintape

⋮

    \endtape
    \bye

The lines marked * can be omitted. All ⟨words⟩ can be omitted, but sometimes have to be replaced by {}, see below.

Comment and empty frames

The frames are: \beginside. . . \endside, \beginfront. . . \endfront and \beginflap. . . \endflap. In each

of these frames you can put normal comment, because each frame is like a miniscule page. If you put a \title or \\ after comment, you have to put a \par (empty line) between them. If you want a frame to be empty, you have to put \empty between \begin and \end.

About formatted input macros

In the figure above you see a lot of things like ‘\title ⟨Title⟩ - ⟨Author⟩.’ This means you have to type ‘\title,’ ⟨space⟩, some words, ‘ - ,’ another word (one) and a ⟨space⟩ (which is usually a return). You could type, for instance:

    \title Pork soda - Primus

If the last part of a line like this contains a space (i.e., if Primus were Mucky Pup), you’d have to put it between braces, so it would look like this:

    \title Pork soda - {Mucky Pup}

If the first part (Pork soda) happens to contain the exact text it should be followed by (‘ - ’ in this example), you have to put it between braces too (so TEX won’t see the first ‘ - ’):

    \title {Pork - soda} - {Mucky Pup}

If something goes wrong here, you’ll get some error message like ‘missing argument,’ ‘use of … doesn’t match it’s definition’ or ‘runaway argument.’

If some part is enclosed by ⟨space⟩s, you have to replace it by an empty group ({}), because two ⟨space⟩s in a row is just the same as one ⟨space⟩. Have a look at page 202 of the TEXbook and the example files to find out more about this.

Size and dimensions of a cover

You can’t have any influence on the appearance of your covers, except for the margins and the size of it. You can create a library of .cas files, in which you define different sizes for tape covers. They define the heights of the contents, front and flap. You can also change the margins, but you need pretty small margins to fit everything on the cover. Have a look at normal.cas, which produces and small.cas, which produces

.

Changing the font

You can change the current fontsize with the commands \small, \normalsize and \large, which work as in LaTEX. You can change the fontstyle with the commands \rm, \it and \em. You can switch to condensed form using \cond. This only works for the normal-sized roman font, as this is the only one which sometimes needs to be condensed. Condensed text looks ugly, but will make sure long titles and author-names can be used. The macros don’t switch to condensed text automaticly, because some people might rather abbrieviate long titles and names.

If you want to change the definition of the fonts, you only need to change the file cascfont.tex. To get the font called cmrcd7, just copy the file cmr7.mf to cmrcd7.mf, edit cmrcd7.mf and change the lines

    font_identifier:="CMR"; font_size 7pt#;
    u#:=15.5/36pt#;      % unit width

to

    font_identifier:="CMRCD"; font_size 7pt#;
    u#:=.75 * 15.5/36pt#;      % unit width

Suggestions and extensions

Use an \empty front so you can draw a logo. Use the flap to enter information like the type of tape used. I made a macro, so I can type \SA90 to get {\large TDK SA90\qquad}.

More about the macro’s

All the frames mentioned above create a \parbox as in LaTEX. You can type normal text in them, as demonstrated in example 2. The commands \begintape, \endcontents, \endfront and \endflap create a horizontal line. This means you don’t have to put them in the order used above. You could make a tapecover with a double frame to put the songs on, by simply repeating the contents part: .